# Group Relative Attention Guidance for Image Editing

### v1

@@ 1 Introduction
Recently, Diffusion Transformer [21] models have once again advanced the field of text-to-image generation [4, 8]. DIT employs a multi-modal attention mechanism (MM-Attention) [8] as its core to progressively inject semantic information from text into noisy latents, ultimately generating high-quality visual outputs through iterative denoising. Unlike UNet-based models [25, 22, 42, 11, 32] that separate cross-attention and self-attention, the unified attention mechanism of DiTs provides a more holistic contextual understanding. This inherent advantage enables it to perform complex image editing even without task-specific fine-tuning [1, 36]. More recently, models such as Kontext [4, 16] and Qwen-Edit [39] further enhance text-driven editing capabilities by continuing training on specialized instruction-editing datasets, demonstrating powerful controllability and generalization.
Figure 2: presents the visualization of the embedding features input to the attention layer, where a significant bias can be observed across different tokens.
@@
Our investigation begins by analyzing the embedding features within each attention layer [13]. We identify a consistent phenomenon: within each layer, feature values concentrate around a shared bias vector. Based on the formulation of MM-Attention, this bias phenomenon can be interpreted as an intrinsic inductive pattern introduced by the architecture itself. We hypothesize that the variation of individual tokens from this bias encodes crucial contextual understanding (the theoretical analysis is presented in Section 4).
This insight directly motivates our method, Group Relative Attention Guidance (GRAG), a guidance mechanism also inspired by the Group Relative Policy Optimization (GRPO) [30] strategy.
@@
As illustrated in Figure 3, GRAG first computes the average Key embedding within each token group to determine a collective editing direction (the common bias vector). Then, a weighting coefficient $\lambda$ is used to modulate each token’s $\Delta$ vector relative to the bias, enhancing those aligned with the editing intent while suppressing conflicting ones. This process leads to more precise and controllable editing outputs. We validate our method on state-of-the-art DIT-based editing models [16, 39, 17]. With a fixed guidance scale, our approach achieves a better trade-off between the editing responsiveness and image consistency, while continuous coefficient adjustment on fixed samples yields smooth and progressive editing outputs (as shown in Figure LABEL:fig:main).
@@ 2 Related Work
2.1 Diffusion Transformers
Aligning textual and visual representations remains a central challenge for transformer-based diffusion models. Early work such as DiT [21, 7, 35] replaced U-Net backbones [25, 22, 26] with transformers and introduced adaptive layer normalization to enable class-conditional generation, but this design limits the ability to achieve denser alignment between textual and visual information. More recent advances, such as MM-DiT [8], address this limitation by introducing a unified token space and bidirectional cross-modal attention, allowing text and image tokens to interact within a shared sequence. Combined with multiple text encoders like CLIP [23] and T5 [24], this design significantly enhances text understanding and enables more accurate and coherent text-guided generation. Recent studies have begun to leverage the contextual modeling capability of DiT for image editing tasks. Kontext [16] adopts the same model architecture as FLUX [4] and is further trained on instruction-based editing datasets. In contrast, Qwen-Edit[39] replaces the T5 encoder with a large vision language model [3, 17] to encode both instructions and reference image information.
@@ 2.2 Text-Driven Image Editing
Early works such as InstructPix2Pix [5] demonstrated that synthetic instruction–response pairs can effectively fine-tune diffusion models for image editing, while training-free methods like Textual Inversion and DreamBooth [9, 28] enabled editing with off-the-shelf generative models [25, 9]. Building on this foundation, subsequent editors—including Emu Edit [31], OmniGen [40], HiDream-E1 [6], and ICEdit [43]—enhanced instruction-driven editing through refined datasets and architectures, while LoRA-based methods [12] introduced task-specific parameter tuning for diffusion transformers. Proprietary multimodal systems such as GPT-4V [20] and Gemini [34], along with platforms like Midjourney [19] and RunwayML [29], have further integrated these advances into end-to-end creative workflows. Kontext[16] extends the FLUX[4] MM-DiT model for editing tasks, leveraging its strong contextual modeling capability to achieve high consistency with reference images. In contrast, models such as Qwen-Edit[39] enhance instruction comprehension through vision language models, enabling more complex and flexible editing operations. Despite progress, instruction-driven image editing still faces two major challenges: (i) striking a balance between editing effectiveness and consistency with the original image, and (ii) achieving precise and continuous control over editing effects. To address these challenges, we investigate the attention-layer representations of DiT and propose Group Relative Attention Guidance (GRAG), which enables precise and controllable editing effects.
@@ 3 Preliminaries
Multi-Modal Diffusion Transformers. The multi-modal diffusion transformer framework, known as multi-modal diffusion transformers (MM-DiT) [21, 8], merges both textual and visual modalities to generate images that align with the semantics of the textual inputs. FLUX incorporates a unified text-image self-attention mechanism, which aligns the multimodal information within each MM-DiT layer. Moreover, FLUX enhances the CLIP [23] text encoder by integrating the T5 [24] encoder, significantly improving its text understanding capabilities.
@@ 4 Bias Vector In The Embedding Vectors
The visualization results of $E$ are shown in Figure 4. In the embedding vector space, each dimension index corresponds to a component, where the dark red regions in Figure 4 indicate positions with larger magnitudes that contribute more to the inner product between different token embeddings. By examining the relationship between RoPE (Rotary Position Embedding [33]) and dimension indices, we observe that text embeddings concentrate in low-frequency components associated with semantics, while image embeddings concentrate in high-frequency components capturing spatial relations. This finding suggests that the two modalities are not fully aligned in the shared embedding space. Furthermore, we investigate the distribution of token embeddings in the vector space. Figure 5 presents the mean vector magnitudes and standard deviations across different attention heads, further revealing the presence of a significant bias vector among tokens in the embedding space.
@@ 5 Group Relative Attention Guidance
Figure 8: Visualization results on training-free image editing method. We update the original first-order inversion in StableFlow with a second-order ODE inversion method[37, 27], referred to as StableFlow++.
@@ 6.1 Experiment Setting
Implementation Details. We validate our proposed method against six image editing baselines. Kontext [16], Step1X-Edit [17] and Qwen-Edit [39] are training-based image editing method. For reproducibility, the random seed is fixed to 42. All experiments are conducted with a batch size of 1 and 24 inference steps. The classifier-free guidance [10] parameter is set following the recommended values for each model, 2.5 for Kontext, 6.0 for Step1X-Edit and 4.0 for Qwen-Edit.
Moreover, GRAG is theoretically applicable to regular MM-DiT-based architecture. Therefore, we select three training-free image editing methods based on Flux.1-Dev[16] T2I models (Flowedit [15], Stableflow [1], Stableflow++) to evaluate the generalization ability of our approach. We provide further discussion in Section 7.
Evaluation. We evaluate our method on PIE [14]. This benchmark covers a diverse range of editing tasks, including object addition/removal, style transfer, and pose modification. For quantitative evaluation, we adopt two complementary perspectives. Following previous works, we adopt LPIPS[41] and SSIM[38] as quantitative metrics to evaluate the content preservation ability in non-edited regions. To assess the alignment between editing results and human preference, we employ the image editing reward model EditScore[18]. EditScore is a reward model fine-tuned on Qwen-2.5VL[2], which measures three aspects: consistency with the original image (Cons), prompt following (PF), and overall edit score (EditScore).
Figure 9: Visualization results of CFG and GRAG under different scales. Compared to CFG, GRAG more effectively regulates the influence of editing instructions on the original image, demonstrating a more accurate and continuous guidance process.
@@ 6.3 Quantitative Analysis
Table 1: Quantitative results on Training-Based method.
@@
Table 2: Ablation analysis of different scales for CFG and GRAG.
@@ 6.4 Ablation Study
Difference with CFG. We compare our approach with the mainstream guidance method, Classifier-Free Guidance (CFG). Unlike CFG, which adjusts the denoising direction during the sampling process, our method directly modulates the editing information within the attention layers. As shown in Table 2 and Figure 10, varying the CFG strength yields few differences. In contrast, GRAG enables precise and continuous control over the editing, producing smooth and consistent adjustments as the editing strength increases, the visual comparsion shown in Figure 9. Such controllability is crucial for customized image editing applications.
Effectiveness Analysis of Group Relative. We analyze the influence of the $\lambda$ and $\delta$ parameters in Eq. 9 on the editing results. Three groups of experiments are conducted: only $\lambda$, only $\delta$, and both $\lambda$ and $\delta$ simultaneously. The qualitative results are shown in Figure 9, while the quantitative results on the PIE benchmark are presented in Table 2 and Figure 10. Adjusting $\lambda$ alone shows no significant impact on the editing results, corresponding to the fluctuating curves in Figure 10, which indicates that tuning $\lambda$ cannot effectively control the editing strength. In contrast, jointly adjusting $\lambda$ and $\delta$ enables a certain degree of controllable editing but fails to achieve continuous precision. Moreover, this simultaneous adjustment often degrades visual fidelity, leading to undesirable artifacts such as the distorted flowers in the second column of the bottom-left sample and the visible artifacts in the first column of the bottom-right sample in Figure 9. Adjusting $\delta$ alone yields the best results, corresponding to the smoothest metric variation in Figure 10 and the most continuous editing transitions in Figure 9.
7 Discussion & Limitation
We further examine the applicability of GRAG to editing methods that employ general MM-Attention architectures involving only text and target image tokens. In these methods, GRAG is applied to the attention layers where source image features are injected. As shown in Figure 8, our approach achieves adjustment of the editing results, indicating that GRAG remains effective in general MM-Attention structures. However, its stability in training-free settings is lower than that in training-based models, as evidenced by the quantitative results in Table 3. We attribute this to the fact that GRAG primarily modulates the cross-attention component in MM-Attention (see Figure 6), whereas in untrained T2I models, source image features are introduced through the edit–edit self-attention branch (Figure 6-b). In such cases, applying GRAG will interfere with existing target image representations.
[…]
Flowedit 0.3758 0.8237 6.8794 5.0531 4.6635
+GRAG 0.3670 0.8312 7.2223 4.8954 4.6697
StableFlow 0.3219 0.9185 8.9309 2.2177 2.4573
+GRAG 0.3292 0.9098 8.8731 2.7429 3.0303
StableFlow++ 0.3691 0.8229 7.3599 5.3926 5.0970
+GRAG 0.3595 0.8316 7.7997 4.8395 4.7251
Table 3: Quantitative results on Training-Free method.
@@ 8 Conclusion
A Pytorch Implementation of GRAG
@@
10
11s_idx, e_idx, bias_scale, delta_scale = 4096, 8192, 1.0, 1.05
12group_bias = img_key[:, s_idx:e_idx, :, :].mean(dim=1)
13img_key[:, s_idx:e_idx, :, :] = bias_scale * group_bias +
@@
15
16# Joint attention computation
17joint_query = torch.cat([txt_query, img_query], dim=1)
18joint_key = torch.cat([txt_key, img_key], dim=1)
19joint_value = torch.cat([txt_value, img_value], dim=1)
20
21joint_hidden_states = dispatch_attention_fn(
@@
29)
B Additional Feature Visualization
B.1 Kontext

### v2

@@ 1 Introduction
Recently, Diffusion Transformer [dit] models have once again advanced the field of text-to-image generation [blackforestlabs2024flux1dev, sd3, chen2023pixart] and image editing [video-editing, video-editing-lora, image-editing, labs2025flux, Flowedit, stableflow, wu2025qwen]. DIT employs a multi-modal attention mechanism [sd3] (MM-Attention) as its core to progressively inject semantic information from text into noisy latents, ultimately generating high-quality visual outputs through iterative denoising. Unlike UNet-based models [ldm, podell2023sdxl, boow-vton, ddpm, DDIM] that separate cross-attention and self-attention, the unified attention mechanism of DiTs provides a more holistic contextual understanding. This inherent advantage enables it to perform complex image editing even without task-specific fine-tuning [stableflow, rf-solver]. More recently, models such as Kontext [blackforestlabs2024flux1dev, labs2025flux] and Qwen-Edit [wu2025qwen] further enhance text-driven editing capabilities by continuing training on specialized instruction-editing datasets, demonstrating powerful controllability and generalization.
Figure 2: The visualization of the embedding features input to the attention layer, where a significant bias can be observed across different tokens.
@@
Our investigation begins with an analysis of the embedding features in each attention layer [jin2025massive]. We identify a consistent phenomenon: within each layer, feature values concentrate around a shared bias vector. Based on the formulation of MM-Attention, this bias phenomenon can be interpreted as an intrinsic inductive pattern introduced by the architecture itself. We hypothesize that the variation of individual tokens from this bias encodes crucial contextual understanding (the theoretical analysis is presented in Section 4).
This insight directly motivates our method, Group Relative Attention Guidance (GRAG), a guidance mechanism also inspired by the Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] strategy.
@@
As illustrated in Figure 3, GRAG first computes the average Key embedding within each token group to determine a collective editing direction (the common bias vector). Then, a weighting coefficient $\lambda$ is used to modulate each token’s $\Delta$ vector relative to the bias, enhancing those aligned with the editing intent while suppressing conflicting ones. This process leads to more precise and controllable editing outputs. We validate our method on state-of-the-art DIT-based editing models [labs2025flux, wu2025qwen, liu2025step1x]. With a fixed guidance scale, our approach achieves a better trade-off between the editing responsiveness and image consistency, while continuous coefficient adjustment on fixed samples yields smooth and progressive editing outputs (as shown in Figure LABEL:fig:main).
@@ 2.1 Text-driven Image Editing
Early works such as InstructPix2Pix [brooks2023instructpix2pix] demonstrated that synthetic instruction–response pairs can effectively fine-tune diffusion models for image editing, while training-free methods like Textual Inversion and DreamBooth [gal2022image, ruiz2023dreambooth] enabled editing with off-the-shelf generative models [ldm, gal2022image]. Building on this foundation, subsequent editors—including Emu Edit [sheynin2024emu], OmniGen [xiao2025omnigen], HiDream-E1 [cai2025hidream], and ICEdit [zhang2025context]—enhanced instruction-driven editing through refined datasets and architectures, while LoRA-based methods [hu2022lora] introduced task-specific parameter tuning for diffusion transformers. Proprietary multimodal systems such as GPT-4V [gpt4v] and Gemini [gemini], along with platforms like Midjourney [midjourney] and RunwayML [runway], have further integrated these advances into end-to-end creative workflows. Kontext[labs2025flux] extends the FLUX.1[blackforestlabs2024flux1dev] model for editing tasks, leveraging its strong contextual modeling capability to achieve high consistency with reference images. In contrast, Qwen-Edit [wu2025qwen] and Step1X-Edit [liu2025step1x] enhance instruction comprehension through vision language models, enabling more complex and flexible editing operations. Despite progress, instruction-driven image editing still faces two challenges: (i) striking a balance between editing effectiveness and consistency with the source image, (ii) achieving precise and continuous control over editing effects.
2.2 Editing Strength Control for Image Editing
Recent progress in controlling editing strength mainly focused on text-to-image [blackforestlabs2024flux1dev, sd3] (T2I) generation. Methods [concept_slider, attr_control, fluxspace] such as ConceptSlier [concept_slider] adjust the influence of specific textual tokens to modulate the corresponding visual concepts, while TACA [TACA] improves text–image alignment by reweighting the attention scores of text tokens in multi-modal attention. Although effective in generative settings, these approaches are not directly applicable to real image editing (TI2I), where both the source image and the editing instruction must be jointly considered. To enable controllable editing on real images, Freeflux [freeflux] employs a gated attention mechanism to regulate the spatial extent of edited regions, and SaaS [saas] modulates instruction strength through attention reweighting, but direct manipulation of attention maps often causes artifacts and limits compatibility with fast attention mechanisms such as Flash Attention [flashattn]. We address these challenges by introducing Group Relative Attention Guidance (GRAG), which analyzes internal attention representations of DiT models and modulates embedding features for precise and continuous control of editing strength with minimal overhead.
@@ 3 Preliminaries
Multi-Modal Diffusion Transformers. The multi-modal diffusion transformer framework, known as multi-modal diffusion transformers (MM-DiT) [dit, sd3, attention], merges both textual and visual modalities to generate images that align with the semantics of the textual inputs. FLUX incorporates a unified text-image self-attention mechanism, which aligns the multimodal information within each MM-DiT layer. Moreover, FLUX enhances the CLIP [clip] text encoder by integrating the T5 [t5] encoder, significantly improving its text understanding capabilities.
@@ 4 Bias Vector In The Embedding Vectors
The visualization results of $E$ are shown in Figure 4. In the embedding vector space, each dimension index corresponds to a component, where the dark red regions in Figure 4 indicate positions with larger magnitudes that contribute more to the inner product between different token embeddings. By examining the relationship between RoPE (Rotary Position Embedding [su2024roformer]) and dimension indices, we observe that text embeddings concentrate in low-frequency components associated with semantics, while image embeddings concentrate in high-frequency components capturing spatial relations. This finding suggests that the two modalities are not fully aligned in the shared embedding space. Furthermore, we investigate the distribution of token embeddings in the vector space. Figure 5 presents the mean vector magnitudes and standard deviations across different attention heads, further revealing the presence of a significant bias vector among tokens in the embedding space.
@@ 5 Group Relative Attention Guidance
Figure 8: Visualization results on training-free image editing method. We update the original first-order inversion in StableFlow with a second-order ODE inversion method[rfsolver, rf-inversion], referred to as StableFlow++.
@@ 6.1 Experiment Setting
Baselines. We apply the fixed GRAG parameters across six image editing models to validate GRAG’s modulation effectiveness. Kontext [labs2025flux], Step1X-Edit [liu2025step1x] and Qwen-Edit [wu2025qwen] are Text-Image to Image (TI2I) editing method. And three training-free Text to Image (T2I) editing methods based on Flux.1-Dev[blackforestlabs2024flux1dev] models (Flowedit [Flowedit], Stableflow [stableflow], Stableflow++). In these methods, GRAG is applied to the attention layers where source image features are injected. For training-based methods, GRAG is applied to all timesteps and layers, while for training-free methods, it is applied only during attention injection. The random seed is fixed to 42. All experiments are conducted with a batch size of 1 and 24 inference steps.
We compare our approach with the mainstream guidance method: Classifier-Free Guidance (CFG) and two attention control methods. Gated Attention [freeflux] (denoted as Attn-gate) controls the editing strength by setting a threshold on the attention correlation values computed between tokens, thereby limiting the number of tokens that respond to the editing instruction. In contrast, Attention Reweighting [saas] (Attn-weight) adjusts the editing strength by reweighting the attention scores produced by the SoftMax layer. All controllable editing strength analysis experiments are conducted based on the Qwen-Edit model.
Evaluation. We evaluate our method on PIE [pnp]. This benchmark covers a diverse range of editing tasks, including object addition/removal, style transfer, and pose modification. For quantitative evaluation, we adopt two complementary perspectives. Following previous works, we adopt LPIPS[LPIPS] and SSIM[SSIM] as quantitative metrics to evaluate the content preservation ability in non-edited regions. To assess the alignment between editing results and human preference, we employ the image editing reward model EditScore[EditScore]. EditScore measures three aspects: consistency with the original image (Cons), prompt following (PF), and overall edit score (EditScore).
Figure 9: Visualization results of different editing strength control methods. Attn-gate controls edits by limiting token activation and cannot enhance existing results (replaced with blanks). Red arrows mark visual artifacts. Overall, adjusting $\delta$ achieves the most continuous and precise editing control.
@@ 6.2 Qualitative Analysis
As shown in Figure 8, our approach achieves adjustment of the editing results, indicating that GRAG remains effective in training-free editing method.
@@
Training-Free
Flowedit 0.3758 0.8237 6.8794 5.0531 4.6635
+GRAG 0.3670 0.8312 7.2223 4.8954 4.6697
StableFlow 0.3219 0.9185 8.9309 2.2177 2.4573
+GRAG 0.3292 0.9098 8.8731 2.7429 3.0303
StableFlow++ 0.3691 0.8229 7.3599 5.3926 5.0970
+GRAG 0.3595 0.8316 7.7997 4.8395 4.7251
Table 1: Quantitative results on different image editing methods.
6.3 Quantitative Analysis
As shown in Table 1-Bottom, GRAG also provides performance gains on training-free editing methods. However, its adaptability is relatively lower compared to training-based methods. We attribute this to the fact that GRAG primarily modulates the cross-attention component in MM-Attention (Figure 6), whereas in untrained T2I models, source image features are introduced through the edit–edit self-attention branch (Figure 6-b). We provide further discussion in the Supplementary Material.
@@ 6.3 Quantitative Analysis
Attn-weight = 0.8 0.3223 0.8940 8.5846 8.4503 7.2808
[…]
Attn-weight = 7.0 0.3055 0.9261 8.8301 8.2159 7.2214
Attn-weight = 10.0 0.2916 0.9484 9.1130 6.7797 5.9555
Attn-weight = 16.0 0.2825 0.9560 9.2361 4.5230 3.9953
[…]
Attn-gate = 0.9 0.3094 0.9031 8.6460 8.1206 7.4004
Attn-gate = 0.7 0.3103 0.9033 8.6655 7.9611 7.2531
Attn-gate = 0.3 0.3296 0.8673 8.1051 7.7184 4.6505
@@
Table 2: Continuity and effectiveness analysis of different editing strength control methods. Appropriate parameters are selected for each method within its effective working range.
@@ 6.4 Ablation Study
Comparison with other guidance method. Although Classifier-Free Guidance (CFG) can influence the generation results by adjusting the denoising process, it provides only limited control over editing strength and often leads to a noticeable degradation in image quality when the guidance scale is small (e.g., CFG = 1). Both attention-based control methods produce visible effects on the edited results; however, Attn-gate exhibits discontinuous control and tends to generate distorted outputs, as shown in the second row on the left of Figure 9. Attn-weight achieves a modulation effect most similar to ours, but it may cause incorrect edits—for example, the cat is mistakenly replaced by a horse in Figure 9—and it still fails to ensure smooth and consistent editing, where the flowers first wither and then become unnaturally vibrant as the editing strength increases. As shown in Table 2 and Figure 10, GRAG enables precise and continuous control over the editing, producing smooth and consistent adjustments as the editing strength increases, the visual comparsion shown in Figure 9. Such controllability is crucial for customized image editing applications.
Effectiveness Analysis of Group Relative. We analyze the influence of the $\lambda$ and $\delta$ parameters in Eq. 9 on the editing results. Three groups of experiments are conducted: only $\lambda$, only $\delta$, and both $\lambda$ and $\delta$ simultaneously. The qualitative results are shown in Figure 9, while the quantitative results on the PIE benchmark are presented in Table 2 and Figure 10. Adjusting $\lambda$ alone shows no significant impact on the editing results, corresponding to the fluctuating curves in Figure 10, which indicates that tuning $\lambda$ cannot effectively control the editing strength. In contrast, jointly adjusting $\lambda$ and $\delta$ enables a certain degree of controllable editing but fails to achieve continuous precision. Moreover, this simultaneous adjustment often degrades visual fidelity, leading to undesirable artifacts such as the distorted flowers and the visible artifacts of the horse sample in Figure 9. Adjusting $\delta$ alone yields the best results, corresponding to the smoothest metric variation in Figure 10 and the most continuous editing transitions in Figure 9.
@@ 7 Conclusion
Group Relative Attention Guidance for Image Editing
Supplementary Material
Contents
Appendix A Pytorch Implementation of GRAG
@@
10s_idx, e_idx, bias_scale, delta_scale = 4096, 8192, 1.0, 1.05
11group_bias = img_key[:, s_idx:e_idx, :, :].mean(dim=1)
12img_key[:, s_idx:e_idx, :, :] = bias_scale * group_bias +
@@
14
15# Joint attention computation
16joint_query = torch.cat([txt_query, img_query], dim=1)
17joint_key = torch.cat([txt_key, img_key], dim=1)
18joint_value = torch.cat([txt_value, img_value], dim=1)
19
20joint_hidden_states = dispatch_attention_fn(
@@
28)
Appendix B Theoretical Analysis of Group Relative Attention Guidance
B.1 Toy Experiment of the GRAG Theoretical Analysis
To further validate the theoretical analysis presented in the main paper and Section 5, we design a controlled toy experiment that isolates the effect of Group Relative Attention Guidance (GRAG) on attention score modulation. The aim is to empirically verify how adjusting the deviation scaling factor $\delta$ and the bias scaling factor $\lambda$ influences the distribution of attention across token groups.
Experimental Setup.
We simplify the MM-Attention mechanism in a minimal setting consisting of a single query $q=1$ and three key tokens.
$A(q,k^{i})=\frac{e^{\langle q,k^{i}\rangle}}{e^{\langle q,k_{s}^{1}\rangle}+e^{\langle q,k_{s}^{2}\rangle}+e^{\langle q,k_{t}\rangle}},$ (S1)
where $k_{s}^{1}=1.9$ and $k_{s}^{2}=1.2$ denote two source-image tokens and $k_{t}=3.4$ denotes the editing-text token. Following the setup in Section 5 of the main paper, two tokens form the source image token group, while the remaining token represents the editing text token. The raw inner-product responses $\langle q,k^{i}\rangle$ of these tokens serve as the unmodulated attention logits. This abstraction preserves the structural essence of text–image cross-attention while allowing us to precisely examine how GRAG affects token-wise attention allocation. Ideally, when increasing editing strength toward higher image–content consistency, the model should maintain responsiveness to the editing instruction while smoothly increasing the contribution of selected source-image tokens.
We compare four representative modulation strategies:
  • Attention Weight - $\gamma$: Directly scaling attention scores after the Softmax layer.
  • GRAG - $\lambda$: Modulating only the bias component while holding the deviation term fixed.
  • GRAG - $\lambda,\delta$: Jointly scaling both the bias component and token-level deviations.
  • GRAG - $\delta$: Modulating only the deviation component while keeping the bias fixed.
Experimental Results.
Figure S1 illustrates this desired behavior using a toy example. The horizontal axis denotes the respective modulation parameter, and the vertical axis shows the post-Softmax attention scores of three token groups: suppressed source-image tokens, enhanced source-image tokens, and editing-text tokens. As seen in subfigures (a)–(c), increasing $\gamma$ or $\lambda$, or jointly tuning $\lambda$ and $\delta$, rapidly suppresses attention to the editing-text tokens, causing the editing instruction to collapse and resulting in artifacts or editing failure despite increased attention to the image tokens.
In contrast, deviation-only modulation via GRAG-$\delta$ preserves stable attention on editing-text tokens while smoothly adjusting the relative emphasis among source-image tokens. This yields the most continuous and controllable editing-strength transition, free of undesirable discontinuities or instruction loss. Overall, GRAG-$\delta$ achieves the desired balance between instruction following and image consistency, providing the most reliable and precise editing-strength control among all tested strategies.
Figure S1: Toy Experiment: Effects of different attention modulation strategies on the attention scores within $M$-Attention. GRAG-$\delta$ enhances or suppresses image tokens while preserving responsiveness to the editing instruction, enabling continuous and precise control over editing strength. The left vertical line marks the lower modification boundary, and the $y$-axis denotes the baseline without any modulation.
B.2 Relationship Between Attention Entropy and GRAG
To quantify how GRAG influences the distribution of attention inside MM-Attention, we measure the attention entropy of each attention map. Given an attention distribution $A^{(l)}\in\mathbb{R}^{N}$ from layer $l$, its entropy is defined as:
$H(A^{(l)})=-\sum_{i=1}^{N}A^{(l)}_{i}\log A^{(l)}_{i}$, (S2)
where lower entropy indicates a more concentrated attention pattern and stronger focus on specific tokens.
We compute attention entropy for all layers of the Qwen-Image-Edit model under different GRAG values. Specifically, for each GRAG scale, we extract the attention maps associated with the source-image tokens and compute their entropies using Equation (S2). The aggregated results are shown in Figure S2(a). As the GRAG value increases, the attention entropy in every layer decreases monotonically, indicating that GRAG strengthens the model’s focus on conditioning information and thereby increases the effective editing strength.
To provide a more intuitive interpretation, we visualize the attention maps corresponding to the same settings. As shown in Figure S2(b), increasing the GRAG value causes the query’s attention to progressively concentrate on the original cat contours. This focused attention corresponds to stronger preservation of source-image structure, leading to improved consistency between the edited output and the reference image.
These results collectively demonstrate that GRAG modulates editing strength by adjusting the concentration of cross-attention, aligning the empirical observations with the theoretical analysis presented in the main paper.
Figure S2: Relationship between attention entropy and GRAG. (a) Increasing the GRAG value on source image tokens leads to a monotonic decrease in attention entropy across layers. (b) The query’s attention over source image tokens becomes progressively more concentrated as the GRAG value increases.
B.3 GRAG of Source Image Tokens
To further understand the distinct roles of the bias and variation components in GRAG, we separately perform continuous adjustments on the text-embedding scaling factors $\lambda$ (bias scaling) and $\delta$ (variation scaling).
$\hat{k}_{\mathrm{s}}^{i}=\lambda\cdot k_{\text{bias}}+\delta\cdot\left(k_{\mathrm{s}}^{i}-k_{\text{bias}}\right)$ (S3)
The results are shown in Fig. S3. When only the bias component is modulated, the overall editing effect remains relatively stable, yet noticeable changes in image quality occur. In contrast, when only the variation component is adjusted, the semantic content of the edited image changes rapidly with increasing editing strength, while the overall image quality remains largely consistent. These observations support our interpretation that the bias component corresponds to the model’s intrinsic editing behavior, whereas the variation component encodes the actual content-specific editing signals.
Figure S3: Effects of the bias and variation components in image-key embeddings.
B.4 GRAG of Editing Text Tokens
We conduct the same decomposition-based analysis on the text tokens to examine the roles of the bias and variation components, with results shown in Fig. S4. Similar to the observations for image tokens, the bias component governs the model’s inherent editing behavior, while the variation component drives the content-specific semantic changes. However, applying GRAG to text tokens reveals a noticeably stronger degree of semantic-level editing control, indicating that modulation of text embeddings provides a more direct and expressive pathway for manipulating editing intensity.
Figure S4: Effects of the bias and variation components in text-key embeddings.
Appendix C Limitation & Discussion
We observe that the effectiveness of GRAG is inherently limited when applied to training-free image editing methods. As illustrated in Fig. S5, training-based TI2I models possess a unified architecture in which the editing instruction and source-image information interact within the same MM-Attention layers, yielding better compatibility with GRAG. In contrast, training-free approaches are built upon T2I models that are not originally designed for editing; they rely on additional inversion procedures and attention injection to approximate the interaction between edited and source-image features. This structural mismatch constrains the impact of GRAG and lead to degraded image quality or ineffective edits (Fig. S6).
Figure S5: Structural differences between training-based and training-free editing methods.
Figure S6: Failure cases of applying GRAG to training-free editing methods.
Appendix D Additional Qualitative Results
Figure S7: Additional qualitative results on existing image editing models.
Figure S8: Additional comparison results with existing editing-strength control methods.
Appendix E Additional Feature Visualization